write standard form of equation where slope of passes through points.
m=2, p(1,3)

Answer by venugopalramana(3286)

write standard form of equation where slope of passes through points.
m=2, p=(X1,Y1)=(1,3)
EQN.IS
Y-Y1=M(X-X1)
Y-3=2(X-1)=2X-2
Y=2X+1